What would be the area of a field 87 rods long and 206 and a half yards wide and what would it cost at 26.5 cents per square foot?
Answer by ankor@dixie-net.com(22740)

What would be the area of a field 87 rods long and 206 and a half yards wide and what would it cost at 26.5 cents per square foot?
:
Convert length and width to feet
:
87 * 16.5 = 1435.5 ft
206.5 * 3 = 619.5 ft
:
1435.5 * 619.5 * .265 = $235,662.45